# Color it, Code it, Cancel it: $k$-local dynamical decoupling from classical additive codes

### v1

@@ I Introduction
Dynamical decoupling (DD), originally inspired by nuclear magnetic resonance [1, 2, 3, 4, 5], has become a powerful technique in quantum information processing. Traditionally, DD has been employed to mitigate imperfections in near-term quantum devices by using sequences of single-qubit operations to isolate quantum systems from environmental noise [6, 7, 8, 9]. Sequences that suppress on-site decoherence to high order in time include for example universal single-qubit protocols [8, 7], XY family [10, 11, 12], concatenated sequences [13, 14], and the Uhrig family [15, 16, 17]. DD has also found applications in quantum sensing, such as frequency-selective spectroscopy [18, 19], and in quantum error correction [20, 21].
Beyond decoherence, DD protocols have been effective in suppressing systematic errors originating from high-order qubit interactions, such as residual crosstalk [22, 23, 24] from two-body couplings, leading to advanced sequences such as WAHUHA [25] and fractal decoupling [26]. However, these methods were designed primarily for small systems with limited connectivity. Recent advances aim to develop time-efficient DD sequences that scale to large systems with arbitrary connectivity graphs [27, 28, 24]. Moreover, while conventional DD sequences typically freeze the system dynamics entirely, making them difficult to adapt for designing specific quantum evolutions, recent developments have proposed DD protocols for quantum simulation, including the engineering of target Hamiltonians using average Hamiltonian theory [29].
@@
In this work, we present a unified framework that leverages graph coloring and coding theory to construct time-efficient DD sequences capable of selectively canceling specific local interactions. Crucially, our approach reveals a connection between DD and classical error-detecting codes, allowing classical additive codes to be translated into scalable DD protocols tailored for general $k$-local Hamiltonians [30]. Our framework also supports finite-width pulses with bounded-strength control [31], lifting the need for idealized instantaneous (bang-bang) operations [7].
We begin by identifying a hidden symmetry in the underlying hypergraph coloring problem, which reduces complexity by grouping qubits into equivalence color classes. The efficient DD sequences for these classes are then systematically constructed using classical additive codes. Our framework generalizes and unifies existing approaches [27, 24, 28, 30, 32] while opening promising avenues for quantum simulation and control, as sketched in Fig. 1. These include universal DD protocols that suppress both decoherence and $k$-local residual interactions, as well as the dynamical design of effective Hamiltonians with tailored interactions. This capability has deep implications for near-term quantum computing and simulations, permitting e.g. detailed studies of dynamical phases of matters such as time crystals. Furthermore, mapping DD sequences to classical detection codes can aid in discovering robust sequences using machine learning [33, 34], by significantly reducing the search space for sequences tailored to specific tasks, and potentially providing an efficient means to generate large datasets of decoupling sequences for training.
@@
This manuscript is organized as follows. In Sec. II, we briefly review bang-bang dynamical decoupling, emphasizing the connection between DD sequences and group theory. In Sec. III, we reinterpret the results of Refs. [27, 24] by uncovering a hidden symmetry structure in colored interaction graphs and extend this perspective to general $k$-local Hamiltonians. In Sec. IV, we establish the correspondence between DD sequences and classical additive codes. Sec. V presents two applications that explicitly demonstrate the power of our framework: (i) a general time-efficient decoupling protocol for two- and three-local interactions tailored to superconducting and spin qubits, and (ii) a DD sequence permitting quantum simulating of the Kitaev honeycomb model, which strongly relies on anisotropic interactions, using only isotropic Heisenberg exchange interactions.
II Decoupling sequence from group theory
@@ II.1 Dynamical decoupling
DD offers a general framework for suppressing unwanted interactions arising from arbitrary Hamiltonians by applying sequences of quantum gates. This approach is traditionally employed in scenarios where a quantum system is weakly coupled to an environment, with the objective that, after the DD sequence, the influence of the environment on the system is effectively canceled. In particular, we consider the Hamiltonian [7]
@@
that couples a qubit system with Hamiltonian $H_{S}$ to an environment $H_{E}$ by the interaction $H_{SE}$. Without loss of generality, we assume that
@@
The DD problem involves designing a control Hamiltonian $H_{c}(t)$ acting solely on the system that suppresses the interactions with the environment. This means that after the applying $H_{c}(t)$, the joint system-environment dynamics are governed by the decoupled Hamiltonian
@@
We restrict ourselves to the analysis of control Hamiltonians comprising sequences that alternate between fast single-qubit gates and free evolution intervals, where all free-evolution intervals have the same duration. These admit a natural group-theoretic description and are especially advantageous for compilation, as they simplify both the implementation and analysis of control protocols. In these cases, the effective Hamiltonian $\overline{H}_{tot}$ can be systematically derived using tools such as Floquet theory and Magnus expansion, assuming that the expansion converges [35].
@@
where $\mathcal{T}$ is the time ordering operator and we set $\hbar=1$. If the control Hamiltonian is cyclic with period $T_{c}$, i.e., $H_{c}(t)=H_{c}(t+T_{c})$, and the initial time is at $t=0$, the stroboscopic time evolution of the system and bath at times $nT_{c}$ with $n\in\mathbb{N}$ is described by [36, 37]
@@
When the frequency $1/T_{c}$ is larger than the largest energy scale of the interaction to be removed [38], the effective Hamiltonian $\overline{H}_{tot}$ is well approximated by the first-order Magnus expansion
@@
Higher-order corrections in the Magnus expansion can be systematically bounded using Lemma 4 of Ref. [38].
In this work, we aim to generalize this concept. Rather than focusing solely on suppressing interactions with the environment, our objective is to design DD sequences that can also selectively suppress certain interactions within the system itself. For example, we either completely freeze the system dynamics (i.e., $\overline{H}_{S}=0$) or we selectively retain only some desired interactions in $H_{S}$, thereby engineering specific target Hamiltonians. This capability is particularly useful for digital-analog quantum simulations [39].
In the next section, we will briefly review the bang-bang control scheme [7, 8], which enables a simple design of DD sequences using group theory. In Appendix A, we also review the bounded-strength control scheme [31], that provides a more realistic model assuming that single-qubit gates have a finite duration. Importantly, we emphasize that the framework we introduce in this work applies to both control schemes.
@@ II.2 Bang-bang control sequence
The bang-bang control scheme [7, 8] assumes that single-qubit pulses are instantaneous. The central idea is to define a DD group $\mathcal{G}=\{g_{i}\}$ [7, 31, 31, 24], which acts on the system Hilbert space via a unitary representation $g\to U_{g}$. The bang-bang DD sequence is then specified by an ordered list of group elements $(g_{1},g_{2},\dots,g_{L})$, with total length of the sequence $L=\lambda|\mathcal{G}|$, where $|\mathcal{G}|$ is the order of the group and $\lambda$ is a positive integer, typically $\lambda=1$ [31]. In general, each group element $g_{i}\in\mathcal{G}$ appears exactly $\lambda$ times in the sequence.
A single cycle of the bang-bang control unitary $U_{c}(t)$ has a total duration $T_{c}$ and is defined as a step function of time divided into $L$ equally-spaced intervals of length $\Delta={T_{c}}/{L}$. During each segment, lasting for a time $\tau\in[0,\Delta)$, the control is a constant unitary operator
@@
This control unitary is abruptly switched from $U_{g_{j-1}}$ to $U_{g_{j}}$, corresponding to an ideal instantaneous control pulse [7, 31].
@@
with free-evolution $U_{\Delta}=\mathcal{T}\text{exp}\left(-i\int_{0}^{\Delta}H_{tot}(\tau)d\tau\right)$.
By Schur’s lemma, the twirling channel $\Pi_{\mathcal{G}}(\cdot)$ acts as a projection and retains only the components of $H_{tot}$ that commute with all elements of the group $\mathcal{G}$. When the map $g\to U_{g}$ defines an irreducible representation of $\mathcal{G}$, the projection takes the form [7, 8, 31, 32]
@@
where $d$ is the dimension of the system Hilbert space. In this special case, the DD procedure completely averages out all nontrivial system operators and the effective system Hamiltonian becomes proportional to the identity. This corresponds to a trivial global phase, meaning that the system dynamics are completely frozen.
When the representation is reducible, the twirling operation projects any operator $O$ onto the multiplicity spaces associated with each irreducible representation. By Maschke’s theorem, the system Hilbert space $\mathcal{H}_{S}$ can be decomposed as
$\mathcal{H}_{S}\cong\bigoplus_{\alpha}V_{\alpha}\otimes\mathbb{C}^{m_{\alpha}},$ (11)
where $\{V_{\alpha}\}$ are inequivalent irreducible representations of dimension $d_{\alpha}$, and $m_{\alpha}$ denotes the multiplicity with which $V_{\alpha}$ appears in the decomposition. By Schur’s lemma, the commutant algebra of $\{U_{g}\}$ takes the form
$\bigoplus I_{d_{\alpha}\times d_{\alpha}}\otimes\text{Mat}_{m_{\alpha}}(\mathbb{C}),$ (12)
where $\text{Mat}_{m_{\alpha}}(\mathbb{C})$ denotes the full algebra of $m_{\alpha}\times m_{\alpha}$ complex matrices acting on the multiplicity space. If the characters $\chi_{\alpha}(g)$ of the irreducible representations $V_{\alpha}$ are known, then the action of the twirling operation is
$\Pi_{\mathcal{G}}(O)\cong\bigoplus_{\alpha}\frac{I_{d_{\alpha}\times d_{\alpha}}}{d_{\alpha}}\otimes\text{Tr}_{V_{\alpha}}(P_{\alpha}OP_{\alpha}),$ (13)
where $P_{\alpha}$ is the orthogonal projector onto the $\alpha$-th isotypic component, given by
$P_{\alpha}=\frac{d_{\alpha}}{|G|}\sum_{g\in\mathcal{G}}\chi_{\alpha}^{*}(g)U_{g}.$ (14)
In the special case where all irreducible representations appear with multiplicity one (i.e., the decomposition is multiplicity-free), the twirling map simplifies to a projection of $O$ onto each irreducible block
$\Pi_{\mathcal{G}}(O)=\sum_{\alpha}P_{\alpha}\frac{\text{Tr}(OP_{\alpha})}{d_{\alpha}}.$ (15)
By selecting a group $\mathcal{G}$ whose action on $\mathcal{H}_{S}$ admits a suitable irreducible decomposition, one can engineer the twirling operation to selectively preserve desired interaction terms in a Hamiltonian. In a later section, we will analyze the commutant algebra of $\{U_{g}\}$ directly through the framework of error detection codes.
@@
In this work, we restrict ourselves to the analysis of $\pi$-pulses generating $X,Y,Z$, and $I$ gates on a system of $N$ qubits. Consequently, we choose the DD group $\mathcal{G}$ to be a subgroup of the $N$-qubit Pauli group modulo phases, i.e. $\mathcal{G}\subseteq\mathcal{P}^{N}/\{\pm i,\pm 1\}$. The corresponding unitary representation maps each group element $g\in\mathcal{G}$ to its associated standard Pauli operator acting on the system Hilbert space.
@@
To illustrate this formalism, we consider, for example, one cycle of the standard single-qubit $\rm XY4$ sequence, defined by the unitary time-evolution operator
$U(T_{c})=XU_{\Delta}YU_{\Delta}XU_{\Delta}YU_{\Delta}\ ,$ (16)
@@
where $U_{\Delta}$ is a single-qubit free evolution operator for a time $\Delta$. This sequence has a DD group $\mathcal{G}$ corresponding to the full single-qubit Pauli group. This can be seen explicitly by expressing the sequence as a product of Pauli operators applied before and after each free evolution period and using Pauli matrix identities ($XY=Z$ and cyclic permutations)
$U(T_{c})\equiv(XU_{\Delta}X)(ZU_{\Delta}Z)(YU_{\Delta}Y)(IU_{\Delta}I)\ .$ (17)
@@
This decomposition shows that each interval is conjugated by a Pauli operator and the full sequence effectively performs a twirl over the single-qubit Pauli group. Consequently, the $\rm XY4$ sequence removes all single-qubit terms, leading to a universal suppression of disorder
@@
In bounded-control schemes [31, 30], the requirement of applying arbitrarily strong and instantaneous control pulses $U_{g_{j+1}}U_{g_{j}}^{\dagger}$ at discrete times $t=j\Delta$ is relaxed. Instead, the control strategy involves smoothly interpolating between successive group elements, implementing a continuous transition from $U_{g_{j}}$ to $U_{g_{j+1}}$ over each subinterval. More details on bounded-control can be found in Appendix A.
III Permutation-invariant decoupling group from graph symmetry

### v2

@@ I Introduction
Dynamical decoupling (DD), originally inspired by nuclear magnetic resonance [Hahn1950, Meiboom1958, maudsley1986modified, brinkmann2016introduction, BRINKMANN2025100191], has become a powerful technique in quantum information processing. Traditionally, DD has been employed to mitigate imperfections in near-term quantum devices by using sequences of single-qubit operations to isolate quantum systems from environmental noise [Vitali1999, viola1998bang, viola1999dynamical, duan1999suppressing]. Sequences that suppress on-site decoherence to high order in time include for example universal single-qubit protocols [viola1999dynamical, viola1998bang], XY family [GULLION1990479, Wang2012, ZhiHui2012], concatenated sequences [West2010, pasini2010optimized], and the Uhrig family [Uhrig2007, wang2011protection, yang2008universality]. DD has also found applications in quantum sensing, such as frequency-selective spectroscopy [szankowski2017environmental, Alvarez2011], and in quantum error correction [paz2013optimally, vezvaee2025demonstrationhighfidelityentangledlogical].
Beyond decoherence, DD protocols have been effective in suppressing systematic errors originating from high-order qubit interactions, such as residual crosstalk [Tripathi2022, qiu2021suppressing, evert2024syncopated] from two-body couplings, leading to advanced sequences such as WAHUHA [WAHUHA1968] and fractal decoupling [agarwal2020dynamical]. However, these methods were designed primarily for small systems with limited connectivity. Recent advances aim to develop time-efficient DD sequences suppressing 2-local interactions for large devices, with sequence length depending only on the topology of the connectivity graphs of the device [brown2024efficient, Coote2025, evert2024syncopated, hickman2025crosstalkrobustdynamicaldecouplingbipartitetopology]. Moreover, while conventional DD sequences typically freeze the system dynamics entirely, making them difficult to adapt for designing specific quantum evolutions, recent developments have proposed DD protocols for quantum simulation, including the engineering of target Hamiltonians using average Hamiltonian theory [Choi2020].
@@
In this work, we present a unified framework that enables the systematic construction of time-efficient and robust DD sequences capable of selectively suppressing any chosen $k$-local interactions, as sketched in Fig. 1. Our framework is built on three key ideas. First, we show that any $k$-local Hamiltonian induces a hypergraph symmetry that partitions physical qubits into equivalence color classes. This symmetry enables the design of DD sequences that suppress $k$-local interactions with sequence lengths determined only by the device topology, rather than by system size. This generalizes prior work [brown2024efficient, Coote2025, evert2024syncopated, hickman2025crosstalkrobustdynamicaldecouplingbipartitetopology], which was limited to 2-local interactions. Second, we exploit the established connection between DD and classical additive codes [Rotteler_2006, Bookatz_2016]. This correspondence allows us to identify high-performance DD sequences directly from code parameters and allow us to design hardware-tailored sequences for both superconducting and spin qubit platforms. Furthermore, it opens paths for using DD as a tool in Hamiltonian engineering [Choi2020, votto2024universal, rajabi2019dynamical]. Finally, we show how to construct DD sequences based on the underlying group structure that are intrinsically robust to pulse imperfections. Our framework naturally supports finite-width pulses with bounded-strength control [Viola2003], eliminating the need for idealized instantaneous (bang–bang) operations [viola1998bang]. Moreover, the mapping between DD sequences and classical detection codes can potentially enables the use of machine-learning approaches [tong2024empirical, rahman2024learning, zhang2025learningsteerquantummanybody] by dramatically reducing the search space of candidate sequences and providing an efficient route to generating large, task-specific training datasets.
@@
The flowchart of the framework proposed in this work is summarized in Fig. 2. The manuscript is organized as follows. In Sec. II, we briefly review bang-bang dynamical decoupling, emphasizing the connection between DD sequences and group theory. In Sec. III, we reinterpret the results of Refs. [brown2024efficient, evert2024syncopated] by uncovering a symmetry structure in colored interaction graphs, and extend this perspective to general $k$-local Hamiltonians. We show that by embedding these symmetries into the group structure of DD sequences, one can construct sequences with lengths depending only on the device topology. In Sec. IV, we introduce the correspondence between DD sequences and classical additive codes, which allows us to characterize the properties of DD sequences directly from code parameters. In Sec. V, we present two applications that explicitly demonstrate the power of our framework: (i) a general time-efficient decoupling protocol for two- and three-local interactions tailored to superconducting and spin qubits, and (ii) a DD sequence permitting quantum simulation of the Kitaev honeycomb model, which strongly relies on anisotropic interactions, using only isotropic Heisenberg exchange interactions. Finally, in Sec. VI, we explain how to construct robust implementations of DD sequences using their group structure, before offering concluding remarks in Sec. VII. While significant room for future improvement remains, our work provides the first systematic framework for constructing scalable DD sequences suitable for near-term quantum devices.
II Decoupling sequences from group theory
In this section, we review the basic of dynamical decoupling theory and its formulation in terms of group theory. This section is organized as follows. In Sec. II.1, we introduce DD from Floquet theory and the first-order Magnus expansion. In Sec. II.2, we formulate dynamical decoupling sequences within a group-theoretic framework.
@@ II.1 Dynamical decoupling
DD offers a general framework for suppressing unwanted interactions arising from arbitrary Hamiltonians by applying sequences of quantum gates. This approach is traditionally employed in scenarios where a quantum system is weakly coupled to an environment, with the objective that, after the DD sequence, the influence of the environment on the system is effectively canceled. In particular, we consider the Hamiltonian [viola1998bang]
@@
that couples a qubit system with Hamiltonian $H_{S}$ to an environment Hamiltonian $H_{E}$ by the interaction $H_{SE}$. Without loss of generality, we assume that
@@
The DD problem involves designing a control Hamiltonian $H_{c}(t)$ acting solely on the system that suppresses the interactions with the environment. This means that after applying $H_{c}(t)$, the joint system-environment dynamics are governed by the decoupled Hamiltonian
@@
We restrict ourselves to the analysis of control Hamiltonians comprising sequences that alternate between fast single-qubit gates and free evolution intervals, where all free-evolution intervals have the same duration. These admit a natural group-theoretic description and are especially advantageous for compilation, as they simplify both the implementation and analysis of control protocols. In these cases, the effective Hamiltonian $\overline{H}_{tot}$ can be systematically derived using tools such as Floquet theory and Magnus expansion, assuming that the expansion converges [Casas_2007].
@@
where $\mathcal{T}$ is the time ordering operator and we set $\hbar=1$. If the control Hamiltonian is cyclic with period $T_{c}$, i.e., $H_{c}(t)=H_{c}(t+T_{c})$, and the initial time is at $t=0$, the stroboscopic time evolution of the system and bath at times $nT_{c}$ with $n\in\mathbb{N}$ is described by [mori2023floquet, iserles2002expansions]
@@
When the frequency $1/T_{c}$ is larger than the largest energy scale [Krodjasteh2008], the effective Hamiltonian $\overline{H}_{tot}$ is well approximated by the first-order Magnus expansion
@@
Higher-order corrections in the Magnus expansion can be systematically bounded using Lemma 4 of Ref. [Krodjasteh2008].
In this work, we aim to generalize this idea. Rather than focusing solely on suppressing interactions with the environment, our objective is to design DD sequences that can also selectively suppress certain interactions within the system itself. For example, we either completely freeze the system dynamics (i.e., $\overline{H}_{S}=0$) or we selectively retain only some desired interactions in $H_{S}$, thereby engineering specific target Hamiltonians. This capability is particularly useful for digital-analog quantum simulations [Parra-Rodriguez2020].
In the next section, we will briefly review the bang-bang control scheme [viola1998bang, viola1999dynamical], which assumes instantaneous single-qubit pulses. In Appendix A, we also review the bounded-strength control scheme [Viola2003], which provides a more realistic model assuming that single-qubit gates have a finite duration. Importantly, we emphasize that the framework we introduce in this work applies to both control schemes.
@@ II.2 Bang-bang control sequence
The bang-bang control scheme [viola1998bang, viola1999dynamical] assumes that single-qubit pulses are instantaneous. The central idea is to define a decoupling group $\mathcal{G}=\{g_{i}\}$ [viola1998bang, Viola2003, Viola2003, evert2024syncopated], which acts on the system Hilbert space via a unitary representation $g\to U_{g}$. The bang-bang DD sequence is then specified by an ordered list of group elements $(g_{1},g_{2},\dots,g_{L})$, with total length of the sequence $L=\lambda|\mathcal{G}|$, where $|\mathcal{G}|$ is the order of the group and $\lambda$ is a positive integer, typically $\lambda=1$ [Viola2003]. In general, each group element $g_{i}\in\mathcal{G}$ appears exactly $\lambda$ times in the sequence. If the group $\mathcal{G}$ is generated by a set of generators $\{\gamma_{1},\dots\gamma_{n}\}$, we write the group as
$\mathcal{G}=\langle\gamma_{1},\dots,\gamma_{n}\rangle$ (7)
to explicitly denote its generating set.
A single cycle of the bang-bang control unitary $U_{c}(t)$ has a total duration $T_{c}$ and is defined as a step function of time, divided into $L$ equally-spaced intervals of length $\Delta={T_{c}}/{L}$. During each segment, lasting for a time $\tau\in[0,\Delta)$, the control is a constant unitary operator
@@
This control unitary is abruptly switched from $U_{g_{j-1}}$ to $U_{g_{j}}$, corresponding to an ideal instantaneous control pulse [viola1998bang, Viola2003].
@@
with the free evolution operator defined as $U_{\Delta}=\mathcal{T}\text{exp}\left(-i\int_{0}^{\Delta}H_{tot}(\tau)d\tau\right)$.
By Schur’s lemma, the twirling channel $\Pi_{\mathcal{G}}(\cdot)$ acts as a projection and retains only the components of $H_{tot}$ that commute with all elements of the group $\mathcal{G}$. When the map $g\to U_{g}$ defines an irreducible representation of $\mathcal{G}$, the projection takes the form [viola1998bang, viola1999dynamical, Viola2003, Rotteler_2006]
@@
where $d$ is the dimension of the system Hilbert space. In this special case, the DD sequence completely averages out all nontrivial system operators and $\bar{H}_{\rm S}$ becomes proportional to the identity. This corresponds to a trivial global phase, meaning that the system dynamics are completely frozen.
@@
In this work, we restrict ourselves to the analysis of $(\pi)$ and/or $(-\pi)$-pulses generating $X,Y,Z$, and $I$ gates on a system of $N$ qubits. Consequently, we choose the decoupling group $\mathcal{G}$ to be a subgroup of the $N$-qubit projective Pauli group, i.e. $\mathcal{G}\subseteq\mathcal{P}^{N}/\{\pm i,\pm 1\}$. We remark that although the Pauli group is not Albelian (i.e $g_{1}g_{2}\neq g_{2}g_{1}$), the projective Pauli group is. The corresponding unitary representation maps each group element $g\in\mathcal{G}$ to its associated standard Pauli operator acting on the system Hilbert space.
@@
As an example, we consider one cycle of the standard single-qubit $\rm XY4$ sequence, defined by the unitary time-evolution operator
$U(T_{c})=U_{\Delta}YU_{\Delta}XU_{\Delta}YU_{\Delta}X\ ,$ (12)
@@
where $U_{\Delta}$ is a single-qubit free evolution operator for a time $\Delta$. This sequence has a decoupling group $\mathcal{G}$ corresponding to the full single-qubit Pauli group. This can be seen explicitly by expressing the sequence as a product of Pauli operators applied before and after each free evolution period and using Pauli matrix identities ($XY=Z$ and cyclic permutations)
$U(T_{c})\equiv(IU_{\Delta}I)(YU_{\Delta}Y)(ZU_{\Delta}Z)(XU_{\Delta}X)\ .$ (13)
@@
This decomposition shows that each interval is conjugated by a Pauli operator and the full sequence effectively performs a twirl over the single-qubit Pauli group. Consequently, the $\rm XY4$ sequence removes all single-qubit terms, resulting in universal noise suppression.
@@
In bounded-control schemes [Viola2003, Bookatz_2016], the requirement of applying arbitrarily strong and instantaneous control pulses $U_{g_{j+1}}U_{g_{j}}^{\dagger}$ at discrete times $t=j\Delta$ is relaxed. Instead, the control strategy involves smoothly interpolating between successive group elements, implementing a continuous transition from $U_{g_{j}}$ to $U_{g_{j+1}}$ over each subinterval. More details on bounded-strength control scheme can be found in Appendix A.
III Permutation-invariant decoupling group from hypergraph symmetry